# On the Failure of Latent State Persistence in Large Language Models

### v4

@@ 1 Introduction
Our framework distinguishes itself from existing “working memory” benchmarks, such as N-back tasks (gong2024working; zhang2024working), which often conflate internal maintenance with context-window retrieval (Appendix A.2). By ensuring the target state is never present in the input, our methodology provides a faithful evaluation of true working memory. Our contributions are as follows:
@@ 2.2 Results
Table 1 reports the ESM across all models. We highlight three critical observations: (1) Non-monotonic Scaling: Model performance does not scale linearly with size or recency. In the GPT family, GPT-4o-0806 yields the highest ESM, outperforming both its successor (GPT-4o-1120) and GPT-4. Notably, LLaMA-3.1-8B achieves a higher ESM than its larger 70B and 405B counterparts. (2) Invariance to Reasoning Traces: Contrary to expectations, models featuring internal CoT—such as o1, o3, and DeepSeek-R1—do not bridge the LSP gap. Despite extended compute, these models still fail to instantiate a continuous and consistent LSP, treating the task as a reactive prompt-completion exercise. (3) Stochasticity of Capability: The acquisition of this state-like behavior appears stochastic rather than a predictable emergent property of scale, suggesting that current training objectives do not explicitly incentivize latent state maintenance.
@@
To verify if these failures are artifacts of decoding, we evaluated the ESM under varying Temperature ($T$) and Top-$p$ settings with GPT-4o-2024-08-06. As shown in Table 3, while higher $T$ increases the variance of responses and lower Top-$p$ concentrates more mass on the number 7, the qualitative conclusion does not change: the ESM of this model never converges to zero as other models show. This confirms that the lack of commitment is not a surface-level decoding issue but a fundamental architectural incapacity to maintain internal variables outside the explicit context.
@@ 3.1 Methodology
An agent possesses exhibiting perfect LSP must instantiate and maintain a fixed latent variable $x$ throughout the trial $\mathcal{T}$. For any query $Q_{t}$ involving a reference object $x_{\text{ref}}\in\mathcal{X}$, an attribute $a$, and a comparison operator $\text{op}\in\{>,<\}$, the response $y_{t}$ is strictly determined by:
@@
Since the inference complexity is low, an agent with perfect LSP can keep all responses logically consistent, i.e., $\mathcal{S}_{t}\neq\emptyset$.
@@ 3.2 Results
Our findings indicate that Hint provides marginal improvement, confirming that contradictions stem from logical drift rather than factual ignorance. Conversely, Explicit Grounding substantially elevates MSC and PR, as it collapses the DCSP into a series of independent, short-context comparisons. This confirms that the primary bottleneck is the internal maintenance of a consistent latent variable $x$.
@@ 4 Mathematical Mentalism
To stress-test the limits of LSP, we design a task inspired by mentalism—a branch of magic that simulates mind-reading through deterministic mathematical routines. Unlike the Kaprekar routine (kaprekar1955interesting), which converges to the constant $1089$ and is likely present in the training corpora of LLMs, we employ a more complex, multi-step sequence based on a variant of the Josephus Problem (schumer2002josephus). This routine involves hidden state initialization followed by a series of permutations and eliminations that yield a predictable invariant, as illustrated in Table 7.
@@ 4.1 Methodology
The routine is designed to satisfy an invariant property $\mathcal{I}$: for any $\mathbf{x}_{0}$ sampled from a valid initialization space, the terminal state $\mathbf{x}_{K}=\mathcal{F}(\mathbf{x}_{0})$ must consist of exactly two identical elements, i.e., $|\mathbf{x}_{K}|=2$ and $x_{K,1}=x_{K,2}$. The divergence between human-like reasoning and autoregressive generation can be formalized through the mechanism of state tracking.
@@
For a deterministic routine $\mathcal{F}$ satisfying $\mathcal{I}$, an agent with perfect LSP can maintain $\mathbf{x}_{t}$ at each $t\in[1,K]$, ensuring the terminal output $x_{K,1}=x_{K,2}$ regardless of the unobserved $\mathbf{x}_{0}$.
Conversely, an agent lacking LSP generates the terminal response $x_{K}$ by marginalizing over the latent trajectory in the linguistic token space:
$P(x_{K}\mid\mathcal{T})=\sum_{\hat{\mathbf{x}}_{0}}P(y\mid\mathcal{F},\hat{\mathbf{x}}_{0})P(\hat{\mathbf{x}}_{0}\mid P_{\text{inst}},R_{\text{ack}}).$ (6)
@@
$\text{ISR}(M)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(x_{K}\text{ satisfies }\mathcal{I}),$ (7)
@@
Table 7: Sequence of deterministic operations for the Mathematical Mentalism task. The variables $a\in\{1,\dots,7\}$ and $b,c\in\{1,\dots,3\}$ help to define the state transitions $f\in\mathcal{O}$.
@@ 4.2 Results
Empirical evaluation across 17 SOTA models reveals a pervasive failure to satisfy the invariant property $\mathcal{I}$ in the absence of explicit state tracking (Table 8). The majority of models exhibit near-zero ISR, suggesting that autoregressive transformers cannot maintain the integrity of a hidden variable $\mathbf{x}_{t}$ through a sequence of non-linear transformations. Notably, the LLaMA model family demonstrates a marginally higher ISR and, as shown in Table 1, generates a more diversified distribution of outcomes. This suggests that while LLaMA’s pre-training may capture more robust mathematical priors, it remains fundamentally bounded by the LSP gap, failing to simulate the deterministic evolution of the state vector $\mathbf{x}$.
@@ 5 Discussion
Appendix A Memory of Humans and LLMs
@@
Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response. [^3] A possibility remains that your response was given by chance if you tell a lie (Figure 4 Case 2) or do not think of a number at all (Figure 4 Case 3).
Memory has attracted increasing attention from both industry and research communities. OpenAI was the first to introduce a memory module in ChatGPT (February 2024) [^4] https://openai.com/index/memory-and-new-controls-for-chatgpt/ that allowed the model to remember information from previous interactions with a user, such as the user’s facts and preferences. By mid-2025, xAI, Anthropic, and Google have integrated memory into Grok, [^5] https://x.com/grok/status/1912670182012801156 Claude, [^6] https://www.anthropic.com/news/claude-4 and Gemini, [^7] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively.
A.1 What Is memory?
atkinson1968human categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime). Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning (baddeley1974working). These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) (weng2023agent). Researchers in the AI community have investigated long-term memory mechanisms for both individual LLMs (wu2025human; du2025rethinking) and LLM agents (zhang2025g; xu2025mem). Recent work has shifted from context-dependent approaches (e.g., Chain-of-Thought (CoT) (wei2022chain) and scratchpads (lanchantin2023learning)) to explicit storage methods (e.g., text-based (park2023generative) or vector-based (hatalis2023memory)) to support lifelong memory (zheng2025lifelong; wang2025memoir). Most studies frame memory as an engineering problem: enabling LLMs to store and retrieve information for later use. By contrast, working memory remains relatively underexplored.
A.2 Evaluation of Working Memory
@@
A.3 Relation between Our Framework and Working Memory
@@
Appendix B Sample Size

### v5

@@ 1 Introduction
Our framework distinguishes itself from existing “working memory” benchmarks, such as N-back tasks (gong2024working; zhang2024working), which often conflate internal maintenance with context-window retrieval (Appendix B.2). By ensuring the target state is never present in the input, our methodology provides a faithful evaluation of true working memory. Our contributions are as follows:
@@ 2.2 Results
Table 1 reports the ESM across all models. We highlight three critical observations: (1) Non-monotonic Scaling: Model performance does not scale linearly with size or recency. In the GPT family, GPT-4o-0806 yields the highest ESM, outperforming both its successor (GPT-4o-1120) and GPT-4. Notably, the ESM of LLaMA-3.1-8B is closer to one, compared to its larger 70B and 405B counterparts. (2) Invariance to Reasoning Traces: Contrary to expectations, models featuring internal CoT—such as o1, o3, and DeepSeek-R1—do not bridge the LSP gap. Despite extended compute, these models still fail to instantiate a continuous and consistent LSP, treating the task as a reactive prompt-completion exercise. (3) Stochasticity of Capability: The acquisition of this state-like behavior appears stochastic rather than a predictable emergent property of scale, suggesting that current training objectives do not explicitly incentivize latent state maintenance.
@@
To verify if these failures are artifacts of decoding, we evaluated the ESM under varying Temperature ($T$) and Top-$p$ settings with GPT-4o-2024-08-06. As shown in Table 3, while higher $T$ increases the variance of responses and lower Top-$p$ concentrates more mass on the number 7, the qualitative conclusion does not change: the ESM of this model never converges to zero as other models show. This confirms that the capacity to maintain internal variables outside the explicit context is not related to surface-level decoding but a more fundamental architectural issue.
@@ 3.1 Methodology
An agent exhibiting perfect LSP must instantiate and maintain a fixed latent variable $x$ throughout the trial $\mathcal{T}$. For any query $Q_{t}$ involving a reference object $x_{\text{ref}}\in\mathcal{X}$, an attribute $a$, and a comparison operator $\text{op}\in\{>,<\}$, the response $y_{t}$ is determined by:
@@
The ability of an agent with perfect LSP to keep all responses logically consistent, i.e., $\mathcal{S}_{t}\neq\emptyset$, is unrelated to $t$.
@@ 3.2 Results
Our findings indicate that Hint even decreases the performance, confirming that contradictions stem from logical drift rather than factual ignorance. Conversely, Hint + All substantially elevates MSC and PR, as it collapses the DCSP into a series of independent, short-context comparisons. This confirms that the primary bottleneck is the internal maintenance of a consistent latent variable $x$.
@@ 4 Mathematical Mentalism
To stress-test the limits of LSP, we design a task inspired by mentalism—a branch of magic that simulates mind-reading through mathematical routines. Unlike the Kaprekar routine (kaprekar1955interesting) and the 1089 trick, [^3] https://en.wikipedia.org/wiki/1089_(number) which converge to a constant (6174 or 1089) and are likely present in the training corpora of LLMs, we employ a more complex, multi-step sequence based on a variant of the Josephus Problem (schumer2002josephus). This routine involves a series of permutations and eliminations that yield a predictable invariant, as illustrated in Table 7.
@@ 4.1 Methodology
The routine is designed to satisfy an invariant property $\mathcal{I}$: for any $\mathbf{x}_{0}$ sampled from a valid initialization space, the terminal state $\mathbf{x}_{K}=\mathcal{F}(\mathbf{x}_{0})$ must consist only one element identical to a hidden one $x_{\text{hidden}}$ in the middle, i.e., $x_{K,1}=x_{\text{hidden}}$. The divergence between human-like reasoning and autoregressive generation can be formalized as the mechanism of state tracking.
@@
For a routine $\mathcal{F}$ satisfying $\mathcal{I}$, an agent with perfect LSP can maintain $\mathbf{x}_{t}$ at each $t\in[1,K]$, ensuring the terminal output $x_{K,1}=x_{\text{hidden}}$ regardless of the unobserved $\mathbf{x}_{0}$.
We prove it in Appendix A. Conversely, an agent lacking LSP generates the terminal response $x_{K}$ by marginalizing over the latent trajectory in the linguistic token space:
$P(x_{K}\mid\mathcal{T})=\sum_{\hat{\mathbf{x}}_{0}}P(x_{K}\mid\mathcal{F},\hat{\mathbf{x}}_{0})P(\hat{\mathbf{x}}_{0}\mid P_{\text{inst}},R_{\text{ack}}).$ (6)
@@
$\text{ISR}(M)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(x_{K,1}=x_{\text{hidden}}),$ (7)
@@
Table 7: Sequence of operations for the Mathematical Mentalism task. The variables $a\in\{1,\dots,7\}$ and $b,c\in\{1,\dots,3\}$ help to define the state transitions $f\in\mathcal{O}$.
@@ 4.2 Results
Empirical evaluation across 17 SOTA models reveals a pervasive failure to satisfy the invariant property $\mathcal{I}$ in the absence of explicit state tracking (Table 8). The majority of models exhibit near-zero ISR, suggesting that autoregressive transformers cannot maintain the integrity of a hidden variable $\mathbf{x}_{t}$ through a sequence of non-linear transformations. Notably, the LLaMA model family demonstrates a marginally higher ISR and, as shown in Table 1, generates a more diversified distribution of outcomes. This suggests that while LLaMA’s pre-training may capture more robust mathematical priors, it remains fundamentally bounded by the LSP gap, failing to simulate the evolution of the state vector $\mathbf{x}$.
@@ 5 Discussion
Appendix A Proof of the State-Agnostic Invariance
We provide a formal proof of correctness for the Proposition 4.1, which is a variation of the Josephus Problem. We demonstrate that the final remaining element $x_{K,1}$ is invariant to the stochastic parameters $\{a,b,c\}$ and is equal to the sequestered element $x_{\text{hidden}}$.
Initialization and Periodicity Construction.
Let $\mathbf{x}_{0}=[x_{0,1},x_{0,2},x_{0,3},x_{0,4}]$ be the initial vector. By concatenating the sequence with itself (Step 2), we define $\mathbf{x}_{1}=[x_{0,1},x_{0,2},x_{0,3},x_{0,4},x_{0,1},x_{0,2},x_{0,3},x_{0,4}]$. Invariant 1 (Periodicity): For any index $i\in\{1,\dots,4\}$, $x_{1,i}=x_{1,i+4}$.
Cyclic Permutation and Boundary Alignment.
In Step 3, a cyclic shift is applied where $a$ is an arbitrary integer. Since the shift is applied to a sequence with period 4 and length 8, the property $x_{2,i}=x_{2,i+4}$ is preserved. Crucially, $x_{2,4}=x_{2,8}$. In Step 4, the first three elements are moved to the middle of the sequence. Since the elements at indices $\{1,2,3\}$ are moved to positions in $\{2,\dots,7\}$, the element originally at index 4 moves to index 1, while the element at index 8 remains at index 8. Thus, for $\mathbf{x}_{3}$, we have $x_{3,1}=x_{2,4}\quad\text{and}\quad x_{3,8}=x_{2,8}$. Result: $x_{3,1}=x_{3,8}$ because $x_{2,4}=x_{2,8}$.
State Hiding and Tail Preservation.
In Step 5, we define the hidden reference:
$x_{\text{hidden}}=x_{3,1}.$ (8)
The remaining sequence $\mathbf{x}_{4}$ is $x_{3,2},\dots,x_{3,8}]$. By the result above, the tail element of $\mathbf{x}_{4}$ is $x_{4,7}=x_{3,8}=x_{\text{hidden}}$. Steps 6 and 7 perform a middle-insertion of the top $b$ elements. As long as the insertion index $j<|\mathbf{x}_{4}|$, the tail element remains unchanged. In Step 8, we remove the first $c$ elements, where $c\in\{1,2,3\}$. The resulting sequence $\mathbf{x}_{6}$ has length n=7cn=7-c. Since removal occurs from the head, the tail element is preserved:
$x_{6,n}=x_{\text{hidden}}.$ (9)
Convergence via Josephus-style Elimination.
Step 9 applies a cyclic shift to $\mathbf{x}_{6}$. The new position of $x_{\text{hidden}}$ in $\mathbf{x}_{7}$ is given by:
$(n-(7\pmod{n}))\pmod{n}$. (10)
This is adjusted for 1-based indexing, where a result of 0 maps to $n$. Step 10 defines a recursive operator $h(\mathbf{x})$:
  1. 1.
    Remove the element at index 2.
  2. 2.
    Shift the head element to the tail.
  3. 3.
    Repeat 6c6-c times.
We evaluate the trajectory for all possible values of $c$:
$c$ $n$ Target Position in $\mathbf{x}_{7}$ Survival Path
1 6 $6-(7\pmod{6})=5$ $1,2,3,4,\mathbf{5},6]\to[3,4,\mathbf{5},6,1]\to[\mathbf{5},6,1,3]\to[1,3,\mathbf{5}]\to[\mathbf{5},1]\to[\mathbf{5}]$
2 5 $5-(7\pmod{5})=3$ $1,2,\mathbf{3},4,5]\to[\mathbf{3},4,5,1]\to[5,1,\mathbf{3}]\to[\mathbf{3},5]\to[\mathbf{3}]$
3 4 $4-(7\pmod{4})=1$ $\mathbf{1},2,3,4]\to[3,4,\mathbf{1}]\to[\mathbf{1},3]\to[\mathbf{1}]$
Conclusion.
In all valid parameter spaces of $c$, the elimination process converges to the singleton set $\{x_{\text{hidden}}\}$. Therefore, $x_{K,1}=x_{\text{hidden}}$. The identity of the final element is invariant to the choices of $a,b,$ and $c$.
Appendix B Memory of Humans and LLMs
@@
Memory has attracted increasing attention from both industry and research communities. OpenAI was the first to introduce a memory module in ChatGPT (February 2024) [^4] https://openai.com/index/memory-and-new-controls-for-chatgpt/ that allowed the model to remember previous interactions with a user, such as the user’s facts and preferences. By mid-2025, xAI, Anthropic, and Google have integrated memory into Grok, [^5] https://x.com/grok/status/1912670182012801156 Claude, [^6] https://www.anthropic.com/news/claude-4 and Gemini, [^7] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively.
Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness, since clear perception is necessary to perform the comparison and provide a response. [^8] A possibility remains that your response was given by chance if you tell a lie (Figure 4 Case 2) or do not think of a number at all (Figure 4 Case 3).
B.1 What Is memory?
atkinson1968human categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime). Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning (baddeley1974working). These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) (weng2023agent).
Researchers in the AI community have investigated long-term memory mechanisms for both individual LLMs (wu2025human; du2025rethinking) and LLM agents (zhang2025g; xu2025mem). Recent work has shifted their focus from context-dependent approaches (e.g., CoT (wei2022chain) and scratchpads (lanchantin2023learning)) to explicit storage methods (e.g., text-based (park2023generative) or vector-based (hatalis2023memory)) to support lifelong memory (zheng2025lifelong; wang2025memoir). Most studies frame memory as an engineering problem: enabling LLMs to store and retrieve information for later use. By contrast, working memory remains relatively underexplored.
B.2 Evaluation of Working Memory
@@
B.3 Relation between Our Framework and Working Memory
@@
Appendix C Sample Size